# Physics-Enhanced Graph Neural Networks For Soft Sensing in Industrial Internet of Things

### v1

@@ Abstract
The Industrial Internet of Things (IIoT) is reshaping manufacturing, industrial processes, and infrastructure management. By fostering new levels of automation, efficiency, and predictive maintenance, IIoT is transforming traditional industries into intelligent, seamlessly interconnected ecosystems. However, achieving highly reliable IIoT can be hindered by factors such as the cost of installing large numbers of sensors, limitations in retrofitting existing systems with sensors, or harsh environmental conditions that may make sensor installation impractical. Soft (virtual) sensing leverages mathematical models to estimate variables from physical sensor data, offering a solution to these challenges. Data-driven and physics-based modeling are the two main methodologies widely used for soft sensing. The choice between these strategies depends on the complexity of the underlying system, with the data-driven approach often being preferred when the physics-based inference models are intricate and present challenges for state estimation. However, conventional deep learning models are typically hindered by their inability to explicitly represent the complex interactions among various sensors. To address this limitation, we adopt Graph Neural Networks (GNNs), renowned for their ability to effectively capture the complex relationships between sensor measurements. In this research, we propose physics-enhanced GNNs, which integrate principles of physics into graph-based methodologies. This is achieved by augmenting additional nodes in the input graph derived from the underlying characteristics of the physical processes. Our evaluation of the proposed methodology on the case study of district heating networks reveals significant improvements over purely data-driven GNNs, even in the presence of noise and parameter inaccuracies.
@@ I Introduction
The Industrial Internet of Things (IIoT) has been rapidly emerging as an important framework, using interconnected sensors, actuators, and related devices not only for automation, but also for interconnectivity between subparts of the network, thereby facilitating decision-making processes within the industrial ecosystem. IIoT goes beyond traditional data collection, focusing instead on streamlining operations and facilitating seamless communication between subcomponents of the network. This paradigm not only drives digital transformation but also facilitates automation and optimizes efficiency across diverse sectors such as water, manufacturing, energy, infrastructure management, and healthcare [1, 2]. However, The reliability of IIoT strongly depends on the quality of data and information gathered through sensor networks integrated into complex systems [3]. Therefore, achieving highly reliable measurements in IIoT may encounter obstacles such as the cost of installing large numbers of sensors, limitations in retrofitting existing systems with sensors, sparse deployment of sensors, malfunctioning sensors, energy constraints associated with deploying dense sensor networks, and harsh environmental conditions that may make sensor installation impractical [4]. As a result, instead of solely relying on expensive hardware solutions to address these obstacles, they can be effectively addressed through computational sensing techniques, thereby unlocking the full potential of intelligent systems facilitated by IIoT.
@@
This study aims to estimate pressure and temperature values (considered as soft sensors) across the entire district heating network using only the existing mass flow rate sensors. These pressure and temperature measurements are critical for ensuring the efficient and safe operation of district heating networks. They are regularly used in monitoring and controlling system parameters, which leads to optimized network efficiency, ensured safety of the infrastructure, and accurate billing of customers based on the actual energy usage. We rely on characteristic parameters of the network’s components, such as pipe diameter, length, roughness, and valve flow coefficient, to infer pressure and temperature. Therefore, we leverage these characteristic parameters to estimate virtual sensors beyond the raw sensor measurements obtained from mass flow rate sensors. By incorporating this additional physics-based information derived from Darcy-Weisbach and heat transfer laws into the input graph as additional nodes, we introduce a physics-enhanced model for soft sensing. It is worth mentioning that graph data augmentation has proven effective for GNNs in downstream tasks [36]. For instance, edge manipulation improves performance in semi-supervised node classification tasks [37]. Moreover, it has been shown that node augmentation based on convex combinations of existing nodes can generally enhance the GNN performance for downstream tasks [38]. However, our proposed physics-enhanced method augments nodes by considering physics-based equations of the network, resulting in a more interpretable approach. Finally, our experimental results demonstrate that using physics-enhanced GNNs can significantly improve performance, even in scenarios where physical sensor readings are susceptible to noise and physical parameters such as pipe roughness are subject to inaccuracy.
@@ II-B Spectral-Based Graph Convolutional Neural Networks
II-B1 Chebyshev Spectral Graph Convolutional Operator
where $\tilde{\mathbf{L}}=2\mathbf{L}/\lambda_{\text{max}}-\mathbf{I_{n}}$.
@@ II-C Spatial-Based Graph Convolutional Neural Networks
In contrast to spectral-based techniques that perform convolutions in the Fourier domain, spatial-based approaches directly establish convolution processes within the graph domain utilizing node information from nearby neighborhoods.
II-C1 Attention Mechanism for Graph Neural Networks
Graph Attention Networks (GATs): The core operation within GAT [43] involves gathering features from neighboring nodes using an attention mechanism in the following way:
where $K$ denotes the number of multi-head attention.
Graph Transformer: The Graph Transformer [45] is another variant of GNNs that leverages attention mechanisms during message passing. Its formulation is expressed as follows:
where $d$ represents the hidden size of each attention head.
@@ III-A Fusing Physics-Based and Data-Driven Methods
In this study, we propose enhancing the input graph to the GNN with additional augmented nodes derived from physics-based equations. This enhancement allows us to improve the accuracy of soft sensor estimation, expanding the information available for constructing data-driven soft sensor models. While several approaches have previously been proposed to incorporate physics-based information into data-driven models, our proposed approach focuses on adding additional relevant variables to the data-driven estimators’ input. Figure 1 illustrates the framework and architecture of our proposed physics-enhanced soft sensing approach. In contrast to purely data-driven approaches, which aim to learn a mapping function solely from mass flow rate sensors ($\dot{m}$) to the temperature and pressure targets $\mathbf{y}$ (i.e., $\dot{m}\rightarrow\mathbf{y}$), our framework takes the graph with mass flow rate measurements as inputs, combined with physical process stem from components characteristics parameters. In the initial stage, these parameters are employed to calculate physics-based information. Subsequently, they are combined with sensors measuring physical mass flow rates in the second step leading to the construction of a larger graph with more nodes, encompassing both physical mass flow rate sensors and physics-based information as illustrated in Figure 2. Our proposed framework is highly adaptable and can be integrated with various physics-based methods and architectures.
@@ IV-B Networks Implementation and Evaluation
We evaluate different types of architecture for graph layers, including ChebyNet, GATv2, and Graph Transformer, as illustrated in Fig. 1. After augmenting physics-based nodes and constructing the graph, we pass the resulting graph through two graph layers to obtain new node representations. Throughout the training process, we use the mean squared error loss function and the Nesterov accelerated Adam (NAdam) optimizer for updating the weights. The hyperparameters were fine-tuned through a grid search over specific intervals, with parameter selection guided by the optimization of the validation loss. More precisely, the grid search spanned the interval $[0,8]$ for the number of heads in the graph attention layer. Additionally, we explored values for the Chebyshev filter order within the range $[1,8]$. The window size underwent investigation within the intervals $\{2,4,8,12,16\}$, while the hidden dimensions of the graph layer were searched within the set $\{8,16,32,64\}$. Moreover, the hidden dimension of the multilayer perceptron (MLP) was explored across $\{32,64,128,256\}$. It was observed that incorporating skip connections between graph layers led to improved performance, aligning with prior research indicating potential oversmoothing issues with an increased number of layers in GNNs [52]. Finally, the outputs from the two graph layers were concatenated and fed into an MLP for virtual sensor estimation. It is worth mentioning that the Rectified Linear Unit (ReLU) activation function is not well-suited for our case, particularly when a unit remains unactivated by any input within the dataset. This scenario poses a challenge for the gradient-based optimization algorithm, as it struggles to update the weights of the inactive unit. To address this issue, we employed the Scaled Exponential Linear Units (SELU) activation function after each layer. SELU enhances the Rectified Linear Unit (ReLU) by adding a slope on the negative half-axis. Table II provides a comprehensive overview of the optimal hyperparameter settings for our experiments.
@@ IV-D Evaluation Metrics
$RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$ (22)
$MAE=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,$ (23)
$Accuracy=1-\frac{\|\mathbf{Y}-\hat{\mathbf{Y}}\|_{F}}{\|\mathbf{Y}\|_{F}},$ (24)
@@ IV-E Results and Discussion
In this section, we compare the performance of the proposed physics-enhanced approach with the purely data-driven approach. It is worth mentioning that we consider both noise-free and noisy scenarios in our evaluation, justifying the robustness of the compared approaches using the aforementioned metrics. The proposed method for virtual sensor estimation is compatible with different neural networks. In our case study, we evaluate MLP, CNN, ChebyNet, GATv2, and Graph Transformer since they possess distinct characteristics. Here, we consider purely data-driven approaches as our baseline, while the physics-enhanced approach serves as our proposed method.
@@
Table III and Table IV present the performance of virtual sensor estimation using both the proposed physics-enhanced approach and the baseline approach (purely data-driven) under ideal (noise-free) and noisy conditions, respectively. Combining physics enhancement with any model yields superior results, with an average improvement of 52% in terms of RMSE. Notably, the physics-enhanced Graph Transformer, in particular, demonstrates significant improvements, with relative reductions in RMSE and MAE 81%, and 68%, respectively, for ideal scenarios. For noisy cases, the physics-enhanced GATv2 shows the highest improvement with a 60% relative reduction in both RMSE and MAE. Additionally, accuracy increases for every physics-enhanced method, further illustrating the superiority of the proposed approach. Among the evaluated architectures, the three physics-enhanced GNNs (ChebyNet, GATv2, and Graph Transformer) consistently outperform other models. Therefore, these architectures offer an efficient solution for soft sensing, making them potent tools for data processing in district heating networks. In summary, the physics-enhanced approach, particularly when used with the Graph Transformer model, exhibits superior performance in noise-free scenarios. However, in noisy scenarios, GATv2 outperforms other models. Enhanced soft sensing performance is also evident when assessing individual virtual sensors. Figs. 6 and 7 illustrate the MAE for each sensor based on both purely data-driven and physics-enhanced graph-based models which have proven to be the best performing models for both noise-free and noisy cases, respectively. It is evident that the performance is significantly improved for almost all sensors in our experiments leading to overall model performance enhancement as illustrated in Fig 8.
Figures 9 and 10 illustrate the estimation performance of three graph-based models for particular pressure (P16) and temperature (T25) sensors in noisy scenarios, respectively. It is evident that, for all GNN models, the physics-enhanced framework consistently delivers superior and more reliable estimations, while the solely data-driven approach demonstrates weaker performance at major points.
@@ IV-F Complexity Analysis
In this subsection, we exclusively focus on the time complexity of the message passing and graph convolution operations within each model. For ChebyNet, GATv2, and Graph Transformer, computing the representation of each node $v_{i}$ involves considering its neighbors, with the sum of neighbors across all nodes precisely equaling the number of edges. Therefore, the general time complexity for these approaches is $\mathcal{O}(|\mathcal{E}|)$. In the case of the physics-enhanced model, augmentation of physics-based nodes ($|\mathcal{V}_{\text{aug}}|$) increases the number of augmented edges ($|\mathcal{E}_{\text{aug}}|$), resulting in linear growth in time complexity expressed as $\mathcal{O}(|\mathcal{E}|+|\mathcal{E}_{\text{aug}}|)$. In the extreme scenario where the adjacency matrix is fully connected (not applicable in our case due to distance thresholding in equation (21))), the complexity exhibits quadratic growth with respect to the number of nodes, denoted as $\mathcal{O}((|\mathcal{V}|+|\mathcal{V}_{\text{aug}}|)^{2})$.

### v2

@@ Abstract
The Industrial Internet of Things (IIoT) is reshaping manufacturing, industrial processes, and infrastructure management. By fostering new levels of automation, efficiency, and predictive maintenance, IIoT is transforming traditional industries into intelligent, seamlessly interconnected ecosystems. However, achieving highly reliable IIoT can be hindered by factors such as the cost of installing large numbers of sensors, limitations in retrofitting existing systems with sensors, or harsh environmental conditions that may make sensor installation impractical. Soft (virtual) sensing leverages mathematical models to estimate variables from physical sensor data, offering a solution to these challenges. Data-driven and physics-based modeling are the two main methodologies widely used for soft sensing. The choice between these strategies depends on the complexity of the underlying system, with the data-driven approach often being preferred when the physics-based inference models are intricate and present challenges for state estimation. However, conventional deep learning models are typically hindered by their inability to explicitly represent the complex interactions among various sensors. To address this limitation, we adopt Graph Neural Networks (GNNs), renowned for their ability to effectively capture the complex relationships between sensor measurements. In this research, we propose physics-enhanced GNNs, which integrate principles of physics into graph-based methodologies. This is achieved by augmenting additional nodes in the input graph derived from the underlying characteristics of the physical processes. Our evaluation of the proposed methodology on the case study of district heating networks reveals significant improvements over purely data-driven GNNs, even in the presence of noise and parameter inaccuracies. Our code and data are available under https://github.com/EPFL-IMOS/PEGNN_SS.
@@ I Introduction
The Industrial Internet of Things (IIoT) has been rapidly emerging as an important framework, using interconnected sensors, actuators, and related devices not only for automation, but also for interconnectivity between subparts of the network, thereby facilitating decision-making processes within the industrial ecosystem. IIoT goes beyond traditional data collection, focusing instead on streamlining operations and facilitating seamless communication between subcomponents of the network. This paradigm not only drives digital transformation but also facilitates automation and optimizes efficiency across diverse sectors such as water, manufacturing, energy, infrastructure management, and healthcare [1, 2]. However, The reliability of IIoT strongly depends on the quality of data and information gathered through sensor networks integrated into complex systems [3]. Therefore, achieving highly reliable measurements in IIoT may encounter obstacles such as the cost of installing large numbers of sensors, limitations in retrofitting existing systems with sensors, sparse deployment of sensors, malfunctioning sensors, energy constraints associated with deploying dense sensor networks, and harsh environmental conditions that may make sensor installation impractical [4]. As a result, instead of solely relying on expensive hardware solutions to address these obstacles, they can be effectively addressed through computational sensing techniques, thereby unlocking the full potential of intelligent systems facilitated by IIoT. For instance, recent studies present frameworks that optimize sensors’ sampling rates, edge servers’ preprocessing modes, edge–cloud resource allocation, and joint sensor activation with mobile charging scheduling. These approaches can significantly reduce system-wide energy consumption [5, 6].
@@
This study aims to estimate pressure and temperature values (considered as soft sensors) across the entire district heating network using only the existing mass flow rate sensors. These pressure and temperature measurements are critical for ensuring the efficient and safe operation of district heating networks. They are regularly used in monitoring and controlling system parameters, which leads to optimized network efficiency, ensured safety of the infrastructure, and accurate billing of customers based on the actual energy usage. We rely on characteristic parameters of the network’s components, such as pipe diameter, length, roughness, and valve flow coefficient, to infer pressure and temperature. Therefore, we leverage these characteristic parameters to estimate virtual sensors beyond the raw sensor measurements obtained from mass flow rate sensors. By incorporating this additional physics-based information derived from Darcy-Weisbach and heat transfer laws into the input graph as additional nodes, we introduce a physics-enhanced model for soft sensing. It is worth mentioning that graph data augmentation has proven effective for GNNs in downstream tasks [38]. For instance, edge manipulation improves performance in semi-supervised node classification tasks [39]. Moreover, it has been shown that node augmentation based on convex combinations of existing nodes can generally enhance the GNN performance for downstream tasks [40]. However, our proposed physics-enhanced method augments nodes by considering physics-based equations of the network, resulting in a more interpretable approach. It is worth mentioning that our proposed method for physics-based graph node augmentation can be applied to other types of networks. However, we chose to focus on district heating networks for several reasons. Firstly, this type of network has not been extensively studied using GNNs, presenting a unique opportunity to apply and evaluate our method in a relatively unexplored area. Additionally, district heating networks are inherently more complex than other pipeline networks, such as water distribution networks, due to the need to monitor both temperature and pressure. Furthermore, district heating networks are well-documented in terms of physics and fluid dynamics, providing well-defined equations that can be effectively integrated into GNN-based approaches.
Finally, our experimental results demonstrate that using physics-enhanced GNNs can significantly improve performance, even in scenarios where physical sensor readings are susceptible to noise and physical parameters such as pipe roughness are subject to inaccuracy.
@@ II-B Spectral-Based Graph Convolutional Neural Networks
Spectral-Based Graph Convolutional Neural Networks assume that the filter $\mathbf{g}_{\theta}$ is a set of learnable parameters. However, the eigen-decomposition required for this approach has a computational complexity of $\mathcal{O}(N^{3})$, which is expensive for large-scale graphs. For this reason, some approaches such as Chebyshev Spectral Graph Convolutional Operator (ChebyNet) [43] and Fourier Graph Operator (FGO) [44] are proposed to reduce the computational complexity.
@@ II-C Spatial-Based Graph Convolutional Neural Networks
In contrast to spectral-based techniques that perform convolutions in the Fourier domain, spatial-based approaches directly establish convolution processes within the graph domain utilizing node information from nearby neighborhoods. Spatial-Based Graph Convolutional Neural Networks use the idea of message passing as follows:
$\mathbf{m}_{\mathcal{N}}^{(k)}(u)=\mathsf{AGG}\left\{\left(\mathsf{MSG}^{(k)}(\mathbf{h}_{u}^{(k)},\mathbf{h}_{v}^{(k)}),\,\forall v\in\mathcal{N}(u)\right)\right\},$ (3)
$\mathbf{h}_{u}^{(k+1)}=\mathsf{UPDATE}^{(k)}\left(\mathbf{h}_{u}^{(k)},\,\mathbf{m}_{\mathcal{N}}^{(k)}(u)\right),$ (4)
where $\mathbf{h}_{u}^{(k)}$ represents the embedding of node $u$ at the $k$-th layer. The functions responsible for updating ($\mathsf{UPDATE}$) and messaging ($\mathsf{MSG}$) can be realized through any differentiable function, such as a multilayer perceptron (MLP). The aggregation operation ($\mathsf{AGG}$) represents a permutation-invariant aggregation operator. Here, $\mathcal{N}(u)$ denotes the set of neighboring nodes of node $u$. To improve the aggregation layer in GNNs, Graph Attention Networks (GATs) [45, 46] and Graph Transformer [47] are proposed to assign attention weights to each neighbor, regulating their influence during the aggregation process.
@@ III-A Fusing Physics-Based and Data-Driven Methods
In this study, we propose enhancing the input graph to the GNN with additional augmented nodes derived from physics-based equations. This enhancement allows us to improve the accuracy of soft sensor estimation, expanding the information available for constructing data-driven soft sensor models. While several approaches have previously been proposed to incorporate physics-based information into data-driven models, our proposed approach focuses on adding additional relevant variables to the data-driven estimators’ input. Figure 1 illustrates the framework and architecture of our proposed physics-enhanced soft sensing approach. In contrast to purely data-driven approaches, which aim to learn a mapping function solely from mass flow rate sensors ($\dot{m}$) to the temperature and pressure targets $\mathbf{y}$ (i.e., $\dot{m}\rightarrow\mathbf{y}$), our framework takes the graph with mass flow rate measurements as inputs, combined with physical process stem from components characteristics parameters. In the initial stage, these parameters are employed to calculate physics-based information. Subsequently, they are combined with sensors measuring physical mass flow rates in the second step leading to the construction of a larger graph with more nodes, encompassing both physical mass flow rate sensors and physics-based information as illustrated in Figure 2. As illustrated in the Figure, the initial graph is comprised solely of mass flow rate sensors. In our proposed approach, we augment the graph by incorporating specific mass flow rates of water within the pipes, along with network component characteristics such as pipe length ($L$), pipe diameter ($D$), and pipe cross-sectional area. Subsequently, we compute the pressure drop ($\Delta P$) and temperature drop ($\Delta T$) in the water as it moves through the pipes, using established fluid dynamics equations. These computed values are then added as additional nodes, thereby enriching the data representation and providing a more detailed and informative model. Our proposed framework is highly adaptable and can be integrated with various physics-based methods and architectures.
@@ IV-B Networks Implementation and Evaluation
We evaluate different types of architecture for graph layers, including ChebyNet, GATv2, FGO, and Graph Transformer, as illustrated in Fig. 1. After augmenting physics-based nodes and constructing the graph, we pass the resulting graph through two graph layers to obtain new node representations. Throughout the training process, we use the mean squared error loss function and the Nesterov accelerated Adam (NAdam) optimizer for updating the weights. The hyperparameters were fine-tuned through a grid search over specific intervals, with parameter selection guided by the optimization of the validation loss. More precisely, the grid search spanned the interval $[0,8]$ for the number of heads in the graph attention layer. Additionally, we explored values for the Chebyshev filter order within the range $[1,8]$. The window size underwent investigation within the intervals $\{2,4,8,12,16\}$, while the hidden dimensions of the graph layer were searched within the set $\{8,16,32,64\}$. Moreover, the hidden dimension of the multilayer perceptron (MLP) was explored across $\{32,64,128,256\}$. It was observed that incorporating skip connections between graph layers led to improved performance, aligning with prior research indicating potential oversmoothing issues with an increased number of layers in GNNs [54]. Finally, the outputs from the two graph layers were concatenated and fed into an MLP for virtual sensor estimation. It is worth mentioning that the Rectified Linear Unit (ReLU) activation function is not well-suited for our case, particularly when a unit remains unactivated by any input within the dataset. This scenario poses a challenge for the gradient-based optimization algorithm, as it struggles to update the weights of the inactive unit. To address this issue, we employed the Scaled Exponential Linear Units (SELU) activation function after each layer. SELU enhances the Rectified Linear Unit (ReLU) by adding a slope on the negative half-axis. Table II provides a comprehensive overview of the optimal hyperparameter settings for our experiments.
@@ IV-D Evaluation Metrics
$\mathsf{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$ (15)
$\mathsf{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,$ (16)
$\mathsf{Accuracy}=1-\frac{\|\mathbf{Y}-\hat{\mathbf{Y}}\|_{F}}{\|\mathbf{Y}\|_{F}},$ (17)
@@ IV-E Results and Discussion
In this section, we compare the performance of the proposed physics-enhanced approach with the purely data-driven approach. It is worth mentioning that we consider both noise-free and noisy scenarios in our evaluation, justifying the robustness of the compared approaches using the aforementioned metrics. The proposed method for virtual sensor estimation is compatible with different neural networks. In our case study, we evaluate MLP, CNN, ChebyNet, GATv2, FGO, and Graph Transformer since they possess distinct characteristics. Here, we consider purely data-driven approaches as our baseline, while the physics-enhanced approach serves as our proposed method.
@@
Table LABEL:ideal and Table LABEL:noisy present the performance of virtual sensor estimation using both the proposed physics-enhanced approach and the baseline approach (purely data-driven) under ideal (noise-free) and noisy conditions, respectively. Combining physics enhancement with any model yields superior results, with an average improvement of 51% in terms of RMSE. Notably, the physics-enhanced Graph Transformer, in particular, demonstrates significant improvements, with relative reductions in RMSE and MAE 81%, and 68%, respectively, for ideal scenarios. For noisy cases, the physics-enhanced FGO shows the highest improvement with a 64% relative reduction in RMSE. Additionally, accuracy increases for every physics-enhanced method, further illustrating the superiority of the proposed approach. Among the evaluated architectures, the four physics-enhanced GNNs (ChebyNet, GATv2, FGO, and Graph Transformer) consistently outperform other models. Therefore, these architectures offer an efficient solution for soft sensing, making them potent tools for data processing in district heating networks. In summary, the physics-enhanced approach, particularly when used with the Graph Transformer model, exhibits superior performance in noise-free scenarios. However, in noisy scenarios, GATv2 outperforms other models. Enhanced soft sensing performance is also evident when assessing individual virtual sensors. Figs. 6 and 7 illustrate the MAE for each sensor based on both purely data-driven and physics-enhanced graph-based models which have proven to be the best performing models for both noise-free and noisy cases, respectively. It is evident that the performance is significantly improved for almost all sensors in our experiments leading to overall model performance enhancement as illustrated in Fig 8.
Figures 9 and 10 illustrate the estimation performance of three graph-based models for particular pressure (P16) and temperature (T25) sensors in noisy scenarios, respectively. It is clear that the physics-enhanced framework outperforms solely data-driven approach across all GNN models, offering more accurate and reliable estimations. While purely data-driven GNN models utilize interactions among sensor measurements and historical data to learn mappings from mass flow rate values to temperature and pressure, they do not incorporate crucial information about the physical process involved. In contrast, physics-enhanced models integrate insights on how water pressure and temperature of water change through the pipe, factoring in pipe characteristics and mass flow rate measurements. This integration of physical knowledge not only enriches the model’s input but also has the potential to enable the understanding of evolving performance degradation (e.g. pipe leakage) within the district heating network—issues that are not explicitly modeled, resulting in significantly lower estimation errors than those seen in purely data-driven models. Furthermore, integrating physical knowledge helps the model to distinguish between normal operational variations and actual signs of degradation. In a district heating network, minor fluctuations in temperature and pressure are expected due to changes in demand and supply conditions. A purely data-driven model might misinterpret these fluctuations as anomalies, whereas a physics-enhanced model can recognize them as normal variations if they align with the expected physical behavior. Conversely, if the model detects a pattern of deviations that do not fit the physical expectations, it can flag this as a potential indicator of an underlying issue, such as a small leak that is gradually getting worse.
@@ IV-F Complexity Analysis
In this subsection, we exclusively focus on the time complexity of the message passing and graph convolution operations within each model. For ChebyNet, GATv2, and Graph Transformer, computing the representation of each node $v_{i}$ involves considering its neighbors, with the sum of neighbors across all nodes precisely equaling the number of edges. Therefore, the general time complexity for these approaches is $\mathcal{O}(|\mathcal{E}|)$. In the case of the physics-enhanced model, augmentation of physics-based nodes ($|\mathcal{V}_{\text{aug}}|$) increases the number of augmented edges ($|\mathcal{E}_{\text{aug}}|$), resulting in linear growth in time complexity expressed as $\mathcal{O}(|\mathcal{E}|+|\mathcal{E}_{\text{aug}}|)$. In the extreme scenario where the adjacency matrix is fully connected (not applicable in our case due to distance thresholding in equation (14))), the complexity exhibits quadratic growth with respect to the number of nodes, denoted as $\mathcal{O}((|\mathcal{V}|+|\mathcal{V}_{\text{aug}}|)^{2})$. For FGO, the computational complexity follows a log-linear relationship with respect to the number of nodes $(|\mathcal{V}^{h}|)$ in the hypervariate graph structure (details are provided in Appendix), denoted as $\mathcal{O}(|\mathcal{V}^{h}|\log{|\mathcal{V}^{h}|})$, whereas for physics-enhanced FGO model, the complexity is expressed as $\mathcal{O}((|\mathcal{V}^{h}|+|\mathcal{V}^{h}_{\text{aug}}|)\log{(|\mathcal{V}^{h}|+|\mathcal{V}^{h}_{\text{aug}}|)})$.